OpenStudy (anonymous):

If you roll a number cube 72 times, how many times do you expect to roll a 6? A 9 B 12 C 11 D 10

OpenStudy (anonymous):

p(6 occur)= 1/6 E(6occur)= 1/6 *72

OpenStudy (anonymous):

WAT!!!!

OpenStudy (anonymous):

12 times

OpenStudy (anonymous):

oh ya thts more of the languagei speak :D

OpenStudy (anonymous):

the probability of getting 6 the first roll is 1/6 so then you would multiply 1/6 by 72 for all those rolls

OpenStudy (anonymous):

OHHHH

OpenStudy (anonymous):

which is what anfar was trying to say

OpenStudy (anonymous):

But u said it in english